On a recent safari, a group of people and elephants contained 100 knees and 100 trunks. If each person took 3 trunks, how many people and how many elephants went on the safari?
Answer by checkley71(8403)

2P+4E=100 & 3P+E=100 OR E=100-3P SUBSTITUTING 100-3P FOR E IN THE FIRST EQUATION WE GET 2P+4(100-3P)=100 OR 2P+400-12P=100 OR -10P=-300 OR P=30
THEREFORE 2*30+4E=100 OR 60+4E=100 OR 4E=40 OR E=10
THERE WERE 30 PEOPLE AND 10 ELEPHANTS ON THE SAFARI.